Circular No. 8604

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ej, 2005ek, 2005el, 2005em Four new apparent supernovae have been discovered on unfiltered CCD

images: 2005ej by T. Boles (Coddenham, England, 0.35-m reflector), and the rest by LOSS/KAIT (via H. Khandrika, D. R. Madison, M. Baek, K. Shimasaki, and W. Li; cf. IAUC 8602).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005ej Sept. 23.89 18^{h}53^{m}38.^{s}48 +66^{o}13′59′′.3 17.4 12′′.1 E, 3′′.4 N
2005ek Sept. 24.53 3 05 48.96 +36 46 10.6 17.5 66′′.5 E, 60′′ S
2005el Sept. 19.48 5 11 48.72 + 5 11 39.4 18.7 39′′ E, 22′′.6 S
2005em Sept. 27.44 3 13 47.71 − 0 14 37.0 18.1 38′′ E, 7′′.9 S

Approximate magnitudes from the respective observers: SN 2005ej, 1991 Aug. 31, [21.0 (Digitized Sky Survey, red); 1992 June 29, [20.5 (DSS, blue); 2005 May 17, [19.5; July 17, [19.5; Sept. 24.812, 17.4. SN 2005ek in UGC 2526, Sept. 18.51, [19; 25.37, 17.3. SN 2005el in NGC 1819, Sept. 14.50, [19; 25.47, 16.0. SN 2005em in IC 307, Sept. 2.43, [19.5; 28.38, 18.0.

COMET C/2005 S1 (SOHO) Further to IAUC 8592, K. Battams reports a fairly bright Kreutz sun-

grazing comet (with T = 2005 Sept. 29.3 TT) on SOHO images (first available position given below); the comet was at mag ∼ 4.2 with a tail ∼ 9′.3 long on Sept. 28.545 UT at ∼ 17.7R_{⊙}.

Comet  2005 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2005 S1 Sept. 27.721 11^{h}44.^{m}8 − 3^{o}39′ C3 HS 2005-S70

CATALOGUE OF COMETARY ORBITS 2005 A new edition of the Catalogue of Cometary Orbits, with entries com-

plete through August, is now available. The 207-page Catalogue sells for $40.00 ($60.00 for airmail delivery outside North America). There is also an e-mail edition (with just the basic orbital information) for $75.00. Orders should be sent to Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A., by FAX to +1-617-495-7231 or by e-mail to mlohmiller@cfa.harvard.edu. Payment can be by credit card or by check made payable to Smithsonian Astrophysical Observatory.

2005 September 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green